OpenStudy (anonymous):

wwhich expression is equivalent to sec^2x-1?

OpenStudy (jdoe0001):

\(\bf \textit{Pythagorean Identities} \\ \quad \\ sin^2(\theta)+cos^2(\theta)=1 \\ \quad \\ 1+cot^2(\theta)=csc^2(\theta) \\ \quad \\ {\color{brown}{ 1+tan^2(\theta)=sec^2(\theta)}}\) what do you think?